OpenStudy (anonymous):

The excess of an angle over its supplement is 20 degrees. Find the angle.

OpenStudy (anonymous):

100 degrees

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

x + y = 180 x - y = 20 Solve for x. You get 100

OpenStudy (anonymous):

but were using a formula that involves 180-n, the equation for a supplememtary angle

OpenStudy (anonymous):

Then 180 - n = x x - n = 20, x = 100